Next Section | Table of Contents


Site 1203

Transit to Site 1203 and Underway Survey

In the early afternoon of 5 July 2001, the JOIDES Resolution departed Yokohama, Japan, and sailed northeasterly to Detroit Seamount. This volcanic edifice, in the northwest corner of the Pacific Basin at the northern tip of the north-northwest–trending Emperor Seamount chain, was the first of four Emperor Seamounts (Detroit, Nintoku, Ojin, and Koko) scheduled to be drilled during Leg 197.

The JOIDES Resolution arrived in the vicinity of Detroit Seamount in the early morning of 11 July and began a ˜14-hr seismic reflection survey to record multicrossings of two proposed sites, HE-3A and HE-3B. These sites had been selected along single-channel analog records gathered in 1988 by the Thomas Washington (Lonsdale et al., 1993). Profiles of good quality were digitally recorded and presented at low exaggeration on EPC line-scan recorders. At 1645 hr on 11 July, the survey was concluded and the seismic equipment retrieved. The thrusters and hydrophones were extended and the vessel situated over the coordinates of proposed Site HE-3A (which became Site 1203) at 1715 hr. The beacon was deployed at 1807 hr.

Hole 1203A

After the vessel settled on location, the corrected precision depth recorder (PDR) depth referenced to the dual elevator stool (DES) was obtained and indicated 2604.4 m. The bottom-hole assembly (BHA) was made up of a 9.875-in rotary core barrel (RCB), a medium-hard formation C-4 bit, a mechanical bit release, a modified head sub, an outer core barrel, a modified top sub, a modified head sub, seven 8.25-in drill collars, a tapered drill collar, six 5.5-in drill pipe sections, and one crossover sub. Hole 1203A was spudded with the RCB at 0145 hr on 11 July. The drill string was drilled ahead with a wash barrel in place to a depth of 70.6 mbsf, when an electrical fault forced coring operations to be prematurely terminated.

Loss of Power

According to the engine room data management system records, at 0259 hr on 12 July, a fault alarm occurred in the power control console for thruster 1. The console housing these control circuits and breaker (Thyrig Bay 1) is located in the forward thyrig room just below the galley in the forward part of the vessel. The electrical fault forced one of the three main generators providing ship's power to drop off-line. As a result of the rapid drop in the supplied power, the power management system compensated by load-shedding noncritical demand. The main power was stabilized when an additional generator was powered up only 4 min after the event.

At 0305 hr, smoke was detected emanating from the forward thyrig room, and at the same time all thrusters were automatically dropped off-line. The vessel began to drift off location. The drill crew was instructed to immediately stop drilling and hang the drill string off at the DES with the 500-ton elevators. The vessel had drifted off location by as much as 200 m (9% of water depth) before positioning control could be restored. To ensure that no damage was inflicted on the BHA or drill string by the excursion off location, the drill string was recovered and the pin and box connections of the BHA components subjected to a magnetic particle inspection. No cracks or bent tubulars were found.

At 1630 hr on 12 July, Hole 1203A was respudded and drilled ahead with a center bit to 300 mbsf at an average rate of penetration (ROP) of 124 m/hr. After the center bit was recovered, the core barrel was dropped and coring initiated. Sediment coring proceeded rapidly to 462.0 mbsf, where basaltic basement was contacted. A total of 162.0 m of nannofossil calcareous ooze and chalk were cored at an average ROP of 58.9 m/hr. Seventy-eight meters of sediment was recovered for an average recovery of 48.2%.

Rotary coring continued 232.7 m into basement (694.7 mbsf) when operations were suspended to change the bit. The average recovery for basement was 55.8%, representing 129.7 m of basalt and volcaniclastic sediment. The average ROP in basement was 3.8 m/hr. The ROP in basement ranged from a lethargic 1.7 m/hr (Core 197-1203A-34R [608.3–612.3 mbsf]) to a frenetic 115.2 m/hr (Core 197-1203A-24R [521.5–531.1 mbsf]). No problems were encountered with hole stability or erratic torque during the first bit run. The bit accumulated 66.9 hr of rotation before being recovered.


The top drive was set back and the drill string pulled back in the hole to 142 mbsf. A free-fall funnel (FFF) was rigged up in the moonpool and launched at 1015 hr on 17 July. The subsea vibration-isolated television (VIT) camera was then deployed and positioned over the hole to observe the condition of the FFF during the extraction of the drill string. The BHA was pulled clear of the seafloor at 1205 hr. As the bit cleared the throat of the FFF, a 1- to 2-m-long squid was observed embracing the outer core barrel.

A new C-7 hard-formation rotary bit was affixed to a fresh mechanical bit release and made up to the BHA. The FFF was reentered at 0008 hr on 18 July. The drill string was advanced to a depth of 655 mbsf. The top drive was picked up and the drill bit positioned at the bottom of Hole 1203A at 694.7 mbsf. Approximately 3.5 m of soft fill was found at the bottom of the hole. Rotary coring resumed in Hole 1203A at 0400 hr on 18 July and advanced through alternating layers of veined and vesicular basalt flows interbedded with volcaniclastic sediment. Coring continued until, at a total depth of 914.6 mbsf, the time allocated for Site 1203 expired. The ROP while coring with the second bit varied from 1.0 m/hr (Core 197-1203A-48R [727.1–732.0 mbsf]) to 16.5 m/hr (Core 197-1203A-64R [866.6–876.2 mbsf]). Core recovery ranged from 1% to 103% (average recovery = 57.4%).

A total of 219.9 m of basement was cored with the second bit at an average ROP of 2.6 m/hr. The second bit acquired 88.3 rotating hours and was still viable when released at the bottom of the hole. The thickness of basement cored was 452.6 mbsf with 255.87 m recovered (recovery = 56.5%) at an average ROP of 3.1 m/hr. The overall results in Hole 1203A were 300 m drilled, 614.6 m cored, and 333.91 m recovered (recovery = 54.3%). The only hole trouble experienced while drilling or coring was when the drill string stuck after picking off bottom to retrieve Core 197-1203A-49R (732.0–741.6 mbsf). After the driller worked the drill string for 30 min with applications of up to 100,000 lb of overpull, the pipe was freed.

The weather on site was, although foggy, mild and under the influence of a large Pacific high. The air temperatures ranged from 6° to 9°C. Seas and swell never exceeded 5 ft and heave ranged from 1 to 5 ft with 2 ft being typical. These agreeable conditions contributed to the good recovery and extended bit life.

Because of a high oil temperature problem due to a misadjustment of the pump controls, the active heave compensator (AHC) was not online until 16 July. The AHC was used continuously while coring from 665.9 to 905.0 mbsf. It was intentionally turned off while cutting the final core (905.0–914.6 mbsf) to see if recovery would be affected. The recovery dropped from 47.5% to 39.5% on the last two cores. The AHC acquired 92 hr of operation on this site.


In preparation for logging, a 30-bbl sepiolite mud flush was circulated in the hole and a wiper trip conducted between 145 and 914.6 mbsf. Following the routine displacement of the borehole with 280 bbl of sepiolite, the bit was released at the bottom of the hole. The Schlumberger equipment was assembled, and the bottom end of the BHA was placed at 202 mbsf.

During the first logging run, the triple combo was lowered to the bottom of the hole at 915 mbsf without difficulty and logging proceeded downhole (dual laterolog and gamma ray). The second pass was conducted uphole at 274 m/hr from 915 mbsf to the base of the pipe. Excellent data quality and repeatability were recorded along the entire section scanned during these two passes.

During the second logging run, the FMS/sonic tool string also succeeded in reaching the bottom of the hole without difficulty and logging proceeded uphole at 274 m/hr. The first pass recorded FMS and sonic data only in the basement section, from 915 to 420 mbsf. The tool string was lowered to the bottom of the hole, and a second pass was conducted in the basement as well as the sedimentary section. Data quality in the basement was excellent. However, in the sedimentary section overlying basement and within some of the volcaniclastic interbeds within the basalt, the borehole was washed, out and, consequently, the FMS data are not usable.

The third logging run was made with the GBM. The first pass was interrupted by an apparent oversaturation of the x-component of the three-axis measured magnetic field. The problem was quickly repaired and the tool lowered; good quality data were recorded downhole and uphole. The fourth logging run was devoted to the magnetic susceptibility tool. This run was canceled when an electronic power supply fault could not be repaired in a timely fashion. The logging operations were secured at 0200 hr on 25 July.

After the logging equipment was rigged down, the drill string was recovered and the BHA was dismantled in preparation for the short trip to Site 1204. The beacon was successfully recalled and retrieved. After the drilling equipment was secured, the vessel departed at 0730 hr on 25 July to run an underway survey over a northern proposed site, HE-3, which is at ODP Site 883.

Next Section | Table of Contents